write the value(s) that make the denominator(s) zero, then solve the equation.
x-6 x+8
--- + 2 = ---
3x x

Answer by stanbon(75887)

write the value(s) that make the denominator(s) zero, then solve the equation.
x-6 x+8
--- + 2 = ---
3x x
----------------
Denominators are zero when x=0
-----------------
Solve:
Multiply thru by 3x to get:
(x-6)+6x = 3(x+8)
7x-6 = 3x+24
4x= 30
x = 7.5
===============
Cheers,
Stan H.